A block with mass m is dragged across a horizontal surface at constant speed v by a force of magnitude Fapp, which acts at an angle of θ above the horizontal surface. The coefficient of kinetic friction between the block and the surface is coefficient of kinetic friction µk. The normal force exerted on the block by the horizontal surface is:


1
Expert's answer
2021-07-19T09:48:18-0400

The normal force exerted on the block by the horizontal surface is:

"N=mg-F_{app}\sin \theta"